### v2

# Polynomial-order oscillations in geometric discrepancy

Abstract
Let $C\subset\mathbb{R}^{2}$ be a convex body, and for a positive integer $N$, let $\mathcal{P}$ be a configuration of $N$ points in $0,1)^{2}$. The discrepancy of $\mathcal{P}$ with respect to $C$ is defined by
$\mathcal{D}(\mathcal{P},\,C)=\sum_{\mathbf{p}\in\mathcal{P}}\sum_{\mathbf{n}\in\mathbb{Z}^{2}}\mathds{1}_{C}(\mathbf{p}+\mathbf{n})-N|C|,$
and one may estimate how $\mathcal{P}$ deviates from uniformity by averaging the latter quantity over a family of sets. When considering quadratic averages over translated and dilated copies of $C$, one gets the homothetic quadratic discrepancy
$\mathcal{D}_{2}(\mathcal{P},\,C)=\int_{0}^{1}\int_{[0,1)^{2}}\left|\mathcal{D}(\mathcal{P},\,\boldsymbol{\tau}+\delta C)\right|^{2}\,{\rm d}\boldsymbol{\tau}\,{\rm d}\delta.$
We investigate the behaviour of the optimal homothetic quadratic discrepancy, that is
$\inf_{\#\mathcal{P}=N}\mathcal{D}_{2}(\mathcal{P},\,C)\quad\text{as}\quad N\to+\infty.$
Beck [BEC88] and Beck and Chen [BC97] showed that the optimal h.q.d. of convex polygons has an order of growth of $\log N$, and more recently, Brandolini and Travaglini [BT22] proved that the optimal h.q.d. of planar convex bodies with a $\mathcal{C}^{2}$ boundary has an order of growth of $N^{1/2}$. We show that, in general, a single order of growth for the optimal h.q.d. need not exist. First, by an implicit geometric construction of $C$, we obtain prescribed oscillations between $\log N$ and $N^{1/2}$. Second, by a subtler design of $\partial C$ and via Fourier-analytic methods, we obtain prescribed polynomial-order oscillations in the range $N^{\alpha}$ with $\alpha\in(2/5,1/2)$. Moreover, we show that the set of planar convex bodies whose optimal h.q.d. does not admit a single order of growth is residual in the (Hausdorff) metric space of planar convex bodies.
1 Introduction
We define the two-dimensional torus as $\mathbb{T}^{2}=\mathbb{R}^{2}/\mathbb{Z}^{2}$, and we identify $0,1)^{2}$ with $\mathbb{T}^{2}$ via the quotient map $\boldsymbol{\tau}\mapsto\boldsymbol{\tau}+\mathbb{Z}^{2}$. Let $C$ be a planar convex body (that is, a compact convex set with non-empty interior), and let $\mathds{1}_{C}$ stand for its characteristic (indicator) function. Consider the periodization functional ${\mathfrak{P}}\colon L^{1}(\mathbb{R}^{2})\to L^{1}(\mathbb{T}^{2})$ defined by
${\mathfrak{P}}\{\mathds{1}_{C}\}(\mathbf{x})=\sum_{\mathbf{n}\in\mathbb{Z}^{2}}\mathds{1}_{C}(\mathbf{x}+\mathbf{n}).$
The theory of irregularities of distribution, often referred to as discrepancy theory, quantifies how point distributions in a space deviate from uniformity (see [MAT10] and [BDP20] for a general overview and applications). In geometric discrepancy theory, historically initiated by Roth’s seminal work [ROT54], one measures such a difference using sets as test functions and compares the number of points that fall within them with the expected value. We introduce the following notions of discrepancy.
Definition 1.1.
Let $C\subset\mathbb{R}^{2}$ be a convex body, and let $\mathcal{P}_{N}$ be a configuration of $N$ points (that is, a collection of $N$ not necessarily distinct points counted with multiplicity) in $\mathbb{T}^{2}$. We define the discrepancy of $\mathcal{P}_{N}$ with respect to $C$ as
$\mathcal{D}(\mathcal{P}_{N},\,C)=\sum_{\mathbf{p}\in\mathcal{P}_{N}}{\mathfrak{P}}\{\mathds{1}_{C}\}(\mathbf{p})-N|C|.$ (1.1)
Let $\boldsymbol{\tau}\in[0,1)^{2}$ be a translation parameter and let $\delta\geq 0$ be a dilation parameter. We define the homothetic quadratic discrepancy (in short, h.q.d.) of $\mathcal{P}_{N}$ with respect to $C$ as
$\mathcal{D}_{2}(\mathcal{P}_{N},\,C)=\int_{0}^{1}\int_{\mathbb{T}^{2}}\left|\mathcal{D}(\mathcal{P}_{N},\,\boldsymbol{\tau}+\delta C)\right|^{2}\,\mathrm{d}\boldsymbol{\tau}\,\mathrm{d}\delta,$ (1.2)
where $\mathrm{d}\boldsymbol{\tau}$ is the normalised Haar measure on $\mathbb{T}^{2}$ and $\mathrm{d}\delta$ is the Lebesgue measure on $[0,1]$.
It is a classical question in the theory of irregularities of distribution to determine the optimal behaviour of the h.q.d. as $N$ (the number of samples per unit square) goes to infinity. In other words, we are concerned with the quantity
$\inf_{\#\mathcal{P}=N}\mathcal{D}_{2}(\mathcal{P},\,C)\quad\text{as}\quad N\to+\infty.$
Surprisingly, the qualitative behaviour of this quantity depends strongly on the geometry of $C$: for instance, when $C$ is the unit square one obtains a logarithmic order of growth, while when $C$ is the unit disk one obtains a polynomial order. We now present sharp estimates for more general bodies, but first, we introduce suitable notation for orders of growth.
Let $f$ and $g$ be two functions defined on the same domain, and let $q$ be an object; we write $f\lesssim_{q}g$ to indicate that there exists a positive value $c$, that may depend on $q$, such that $f\leq cg$ in the whole domain. If we omit $q$ in the subscript, then $c$ is a positive constant. If the $\lesssim_{q}$ (or $\lesssim$) holds in both senses, then we write $f\approx_{q}g$ (or $f\approx g$).
First, we state a result on the optimal h.q.d. with respect to polygons. Throughout the paper, we assume that $N$ is an integer greater than $1$.
Theorem 1.2 (Beck-Chen).
Let $C$ be a convex polygon. Then, it holds
$\inf_{\#\mathcal{P}=N}\mathcal{D}_{2}(\mathcal{P},\,C)\approx_{C}\log N.$
In particular, the lower bound is due to a classical work of Beck [BEC88], while the upper bound follows from a subsequent work of Beck and Chen [BC97].
In contrast with the latter result, one has the following result on the optimal h.q.d. for convex bodies with sufficiently regular boundaries.
Theorem 1.3 (Brandolini-Travaglini).
Let $C$ be a planar convex body. If the boundary of $C$ is $\mathcal{C}^{2}$ (regardless of curvature), then it holds
$\inf_{\#\mathcal{P}=N}\mathcal{D}_{2}(\mathcal{P},\,C)\approx_{C}N^{1/2}.$
The upper bound is well-known and can be easily shown via a probabilistic argument employing a randomly shifted lattice (for example, see [BCC+19]). On the other hand, the lower bound is due to a recent work of Brandolini and Travaglini [BT22] and requires subtle Fourier-analytic techniques; these provide a starting point for our work. In the same paper, the following result is shown, notably recovering a polynomial order of growth across an interval of exponents.
Theorem 1.4 (Brandolini-Travaglini).
Let $\alpha\in[2/5,1/2)$. Then, there exists a planar convex body $C_{\alpha}$ such that
$\inf_{\#\mathcal{P}=N}\mathcal{D}_{2}(\mathcal{P},\,C_{\alpha})\approx_{\alpha}N^{\alpha}.$
In the current paper, we aim to show that, in general, the optimal h.q.d. does not need to exhibit a single order of growth, and prescribed stationary orders of growth can be achieved. For this purpose, we present two general methods for constructing special sets whose optimal h.q.d. continuously switch between different orders of growth (logarithmic or polynomial).
Our first method employs a simple geometric argument. This allows prescribed oscillations from a logarithmic order of $\log N$ to a polynomial order of $N^{1/2}$, and vice versa. We state our result as follows.
Theorem 1.5.
Let $\left\{\alpha_{i}\right\}_{i\in\mathbb{N}}$ be in $\{0,1/2\}$, let $\left\{\varepsilon_{i}\right\}_{i\in\mathbb{N}}$ be in $(0,\varepsilon]$ for some small positive $\varepsilon$, and let $\left\{q_{i}\right\}_{i\in\mathbb{N}}$ be positive integers. Then there exists a planar convex body $C$ and an increasing sequence $\left\{N_{i}\right\}_{i\in\mathbb{N}}\subset\mathbb{N}$ such that, for every $i\in\mathbb{N}$ such that $\alpha_{i}=0$, it holds
$\log^{1-\varepsilon_{i}}N\leq\inf_{\#\mathcal{P}=N}\mathcal{D}_{2}(\mathcal{P},\,C)\leq\log^{1+\varepsilon_{i}}N\quad\text{for every}\quad N\in[N_{i},N_{i}+q_{i}],$
while, for every $i\in\mathbb{N}$ such that $\alpha_{i}=1/2$, it holds
$N^{1/2-\varepsilon_{i}}\leq\inf_{\#\mathcal{P}=N}\mathcal{D}_{2}(\mathcal{P},\,C)\leq N^{1/2+\varepsilon_{i}}\quad\text{for every}\quad N\in[N_{i},N_{i}+q_{i}].$
This first method obtains the planar convex body $C$ as the limit of an increasing sequence of planar convex bodies; thus, $C$ is constructed implicitly. The simplicity of our first method comes at the cost of limited knowledge on the geometry of the boundary of $C$. In particular, we rely on Lemma 2.1, showing that two bodies with a small symmetric difference have close discrepancies. Further, by pairing the just-mentioned lemma with a classical topological argument, we show that the set of planar convex bodies whose optimal h.q.d. has a single order of growth is meagre in the space of planar convex bodies endowed with the Hausdorff metric.
On the other hand, our second method provides a direct construction. This employs harmonic analysis techniques and subtle geometric estimates on the Fourier transform of sets. We obtain prescribed polynomial-order oscillations in the range $N^{\alpha}$ with $\alpha\in(2/5,1/2)$, and state our second result as follows.
Theorem 1.6.
Let $\left\{\alpha_{i}\right\}_{i\in\mathbb{N}}$ be in $(2/5,1/2)$, let $\left\{\varepsilon_{i}\right\}_{i\in\mathbb{N}}$ be in $(0,\varepsilon]$ for some small positive $\varepsilon$, and let $\left\{q_{i}\right\}_{i\in\mathbb{N}}$ be positive integers. Then there exists a planar convex body $C$ and an increasing sequence $\left\{N_{i}\right\}_{i\in\mathbb{N}}\subset\mathbb{N}$ such that, for every $i\in\mathbb{N}$, it holds
$N^{\alpha_{i}-\varepsilon_{i}}\leq\inf_{\#\mathcal{P}=N}\mathcal{D}_{2}(\mathcal{P},\,C)\leq N^{\alpha_{i}+\varepsilon_{i}}\quad\text{for every}\quad N\in[N_{i},N_{i}+q_{i}].$
The second method involves constructing the boundary of $C$ “by hand”; as we will see, it suffices to design the boundary only in a small neighbourhood of a point, while leaving the remaining part as $\mathcal{C}^{2}$. Moreover, we note that we can alternate exponents within the same range as in Theorem 1.4.
To the best of our knowledge, these two results settle the first examples of sets for which the optimal h.q.d. has prescribed stationary orders of growth. Interestingly, if one further considers full rotations in the averaging process (that is, averaging over the whole similarity group), it is a classical result, obtained independently by Beck [BEC87] and Montgomery [MON94, Ch. 6], that one always obtains an order of growth of $N^{1/2}$. The reader may consult a recent work by Gennaioli and the author [BG24] for further details on the interplay between discrepancy and geometric measure theory.
In the pages that follow, Section 2 contains the implicit geometric construction that leads to Theorem 1.5, while Section 3 contains the direct construction that proves Theorem 1.6. In particular, the proof of Theorem 1.6 requires different techniques to establish lower and upper bounds, and therefore, these are treated in distinct subsections.
2 The first method
Recalling the hypotheses of Theorem 1.5, we have that $\left\{\alpha_{i}\right\}_{i\in\mathbb{N}}$ is a sequence of exponents with values in $\{0,1/2\}$, $\left\{\varepsilon_{i}\right\}_{i\in\mathbb{N}}$ is a (possibly decreasing) sequence in $(0,\varepsilon]$ for some small positive $\varepsilon$, and $\left\{q_{i}\right\}_{i\in\mathbb{N}}$ is a (possibly increasing) sequence of positive integers. We remark that $\alpha_{i}$ prescribes the target order of growth, while $\varepsilon_{i}$ and $q_{i}$ quantify, respectively, how closely and for how long we approximate it.
We begin by presenting a simple but essential auxiliary result.
Lemma 2.1.
Let $\mathcal{P}$ be a configuration of $N$ points in $\mathbb{T}^{2}$, and let $A$ and $B$ be two planar sets such that $\rm{diam}(A)\leq 1$, $A\supset B$, and $|A\setminus B|\leq\eta\leq 1$. Then, it holds
$\left|\mathcal{D}_{2}(\mathcal{P},A)-\mathcal{D}_{2}(\mathcal{P},B)\right|\leq 4N^{2}\eta^{1/2}.$
Proof.
By the reverse triangle inequality, it holds
$\left|\mathcal{D}_{2}(\mathcal{P},A)-\mathcal{D}_{2}(\mathcal{P},B)\right|^{2}\leq 4N^{2}\int_{0}^{1}\int_{\mathbb{T}^{2}}\left|\mathcal{D}(\mathcal{P},\boldsymbol{\tau}+\delta A)-\mathcal{D}(\mathcal{P},\boldsymbol{\tau}+\delta B)\right|^{2}\,\mathrm{d}\boldsymbol{\tau}\,\mathrm{d}\delta.$ (2.1)
Now, since $A\supset B$, it holds
$\mathcal{D}(\mathcal{P},\boldsymbol{\tau}+\delta A)-\mathcal{D}(\mathcal{P},\boldsymbol{\tau}+\delta B)=\mathcal{D}\left(\mathcal{P},\boldsymbol{\tau}+\delta(A\setminus B)\right).$ (2.2)
Since $\rm{diam}(A)\leq 1$, then, for any $\mathbf{p}\in\mathbb{T}^{2}$, the set of $\boldsymbol{\tau}$’s such that
$\mathds{1}_{\boldsymbol{\tau}+\delta(A\setminus B)}(\mathbf{p})\neq 0$
has measure $\delta^{2}|A\setminus B|\leq\delta^{2}\eta$. Therefore, it holds
$\int_{\mathbb{T}^{2}}\left|\mathds{1}_{\boldsymbol{\tau}+\delta(A\setminus B)}(\mathbf{p})\right|^{2}\,\mathrm{d}\boldsymbol{\tau}\leq\delta^{2}\eta.$
Now, by applying the Cauchy–Schwarz inequality, we obtain
$\begin{split}\int_{0}^{1}\int_{\mathbb{T}^{2}}\left|\mathcal{D}\left(\mathcal{P},\,\boldsymbol{\tau}+\delta(A\setminus B)\right)\right|^{2}\,\mathrm{d}\boldsymbol{\tau}\,\mathrm{d}\delta&\leq 2\int_{0}^{1}\int_{\mathbb{T}^{2}}\left|\sum_{\mathbf{p}\in\mathcal{P}}\mathds{1}_{\boldsymbol{\tau}+\delta(A\setminus B)}(\mathbf{p})\right|^{2}\,\mathrm{d}\boldsymbol{\tau}\,\mathrm{d}\delta+2N^{2}\eta^{2}\\ &\leq 2N\sum_{\mathbf{p}\in\mathcal{P}}\int_{0}^{1}\int_{\mathbb{T}^{2}}\left|\mathds{1}_{\boldsymbol{\tau}+\delta(A\setminus B)}(\mathbf{p})\right|^{2}\,\mathrm{d}\boldsymbol{\tau}\,\mathrm{d}\delta+2N^{2}\eta^{2}\\ &\leq 2N^{2}\eta+2N^{2}\eta^{2}\leq 4N^{2}\eta.\end{split}$
Combining the latter with (2.1) and (2.2) yields the claim. ∎
The latter result is all we need to proceed with the proof of Theorem 1.5. We start by considering a sequence of convex planar bodies $\left\{C_{i}\right\}_{i\in\mathbb{N}}$ to be chosen later, but such that each of their diameters is less than $1$ and
[…]
Now, let $\left\{\eta_{i}\right\}_{i\in\mathbb{N}}$ be a (decreasing) sequence with values in $(0,1]$, but to be chosen later as well. Moreover, for every $i\in\mathbb{N}$, we further impose that $C_{i}\subset C_{i+1}$ and that $|C_{i+1}\setminus C_{i}|\leq\eta_{i+1}$.
Figure 1: The construction in the proof of Theorem 1.5.
For the sake of notation, consider the function
$H_{i}(N)=\begin{cases}\log N&\text{if}\quad\alpha_{i}=0\\ N^{1/2}&\text{if}\quad\alpha_{i}=1/2\end{cases}.$
We recall that, by Theorem 1.2 and Theorem 1.3, for every $i\in\mathbb{N}$, it holds
$\inf_{\#\mathcal{P}=N}\mathcal{D}_{2}(\mathcal{P},\,C_{i})\approx_{C_{i}}H_{i}(N).$ (2.3)
We now construct the sequence $\left\{\eta_{i}\right\}_{i\in\mathbb{N}}$ recursively. Suppose $\left\{\eta_{j}\right\}_{j=1}^{i}$ and $\left\{C_{j}\right\}_{j=1}^{i}$ have already been chosen. Then, let $N_{i}$ be such that
$H_{i}^{1-\varepsilon_{i}}(N)\leq\inf_{\#\mathcal{P}=N}\mathcal{D}_{2}(\mathcal{P},\,C_{i})\leq H_{i}^{1+\varepsilon_{i}}(N)\quad\text{for every}\quad N\in[N_{i},N_{i}+q_{i}].$ (2.4)
In particular, notice that such an $N_{i}$ exists, or else it would contradict (2.3). Then, choose $\eta_{i+1}$ in such a way that $\eta_{i+1}\leq\eta_{i}/2$ and
$\eta_{i+1}<(N_{i}+q_{i})^{-4}/16.$
By choosing the whole sequence $\left\{\eta_{i}\right\}_{i\in\mathbb{N}}$ in such a fashion, it follows that the nested sequence of the planar convex bodies $\left\{C_{i}\right\}_{i\in\mathbb{N}}$ converges in the Hausdorff metric to a planar convex body $C$ such that, for every $i\in\mathbb{N}$, it holds
$|C\setminus C_{i}|\leq\sum_{j=i+1}^{\infty}\eta_{j}\leq 2\eta_{i+1}.$
Therefore, by applying Lemma 2.1, we obtain that, for every $i\in\mathbb{N}$, for every $N\in[N_{i},N_{i}+q_{i}]$, and for every configuration $\mathcal{P}$ of $N$ points in $\mathbb{T}^{2}$, it holds
$|\mathcal{D}_{2}(\mathcal{P},\,C)-\mathcal{D}_{2}(\mathcal{P},\,C_{i})|\leq 4N^{2}\eta_{i+1}^{1/2}<1.$
Finally, by pairing the latter inequality with (2.4), the claim of Theorem 1.5 follows.
2.1 The discrepancy of most convex bodies
What follows is inspired by [BRT98, Sec. 4], where the authors applied an analogous argument to study the average decay of the Fourier transform of convex sets. We let $\mathfrak{C}$ be the space of convex bodies in $\mathbb{R}^{2}$ endowed with the Hausdorff metric $d_{H}$ defined by
$d_{H}(A,B)=\max\Big\{\sup_{x\in A}\inf_{y\in B}|x-y|,\,\sup_{y\in B}\inf_{x\in A}|x-y|\Big\}\quad\text{for }A,B\in\mathfrak{C}.$
By the Blaschke selection theorem, the metric space $(\mathfrak{C},d_{H})$ is locally compact, and by the Baire category theorem, it follows that $(\mathfrak{C},d_{H})$ is a Baire space and, in particular, it is nonmeagre.
In [SW65] it is shown that the Hausdorff metric and the symmetric-difference metric $d_{\Delta}(A,B)=|A\Delta B|$ induce the same topology on $\mathfrak{C}$. It is not difficult to adapt the proof of Lemma 2.1 and show that, for a fixed $N\in\mathbb{N}$, the functional
$\mathcal{D}_{2}(\mathcal{P},\,\cdot)\colon(\mathfrak{C},d_{\Delta})\to(0,+\infty)$
is (locally) uniformly continuous in its argument with respect to all configurations $\mathcal{P}$ of $N$ points. Indeed, it is enough to pair the equality $A\Delta B=(A\setminus B)\sqcup(B\setminus A)$ with the fact that $\mathcal{D}(\mathcal{P},\,\cdot)$ is additive on disjoint sets, and note that the condition on the diameter in the hypothesis of Lemma 2.1 is (locally) not restrictive. Hence, we infer that, for a fixed $N\in\mathbb{N}$, the functional
$\inf_{\#\mathcal{P}=N}\mathcal{D}_{2}(\mathcal{P},\,\cdot)\colon(\mathfrak{C},d_{\Delta})\to(0,+\infty)$
is continuous. It is time to state a technical result of Gruber [GRU83].
Lemma 2.2 (Gruber).
Let $\{a_{j}\}_{j\in\mathbb{N}}\subset(0,+\infty)$, let $T$ be a nonmeagre topological space, and let $\{\phi_{j}\}_{j\in\mathbb{N}}$ be a sequence of continuous functionals with $\phi_{j}\,\colon\,T\to(0,+\infty)$. Define
$\mathcal{A}$ $=\big\{x\in T\;\colon\;\phi_{j}(x)=o(a_{j})\,\textnormal{ as }\,j\to+\infty\big\},$
$\mathcal{B}$ $=\big\{x\in T\;\colon\;a_{j}=o(\phi_{j}(x))\,\textnormal{ as }\,j\to+\infty\big\}.$
Then, the following holds:
  1. i)
    If $\mathcal{A}$ is dense in $T$, then, for every $x\in T$ except for a meagre subset, the inequality $\phi_{j}(x)<a_{j}$ holds for infinitely many values of $j$.
  2. ii)
    If $\mathcal{B}$ is dense in $T$, then, for every $x\in T$ except for a meagre subset, the inequality $\phi_{j}(x)>a_{j}$ holds for infinitely many values of $j$.
Finally, by pairing Theorem 1.2 and Theorem 1.3 with the latter lemma and the fact that $\inf_{\#\mathcal{P}=N}\mathcal{D}_{2}(\mathcal{P},\,\cdot)$ is continuous in $(\mathfrak{C},d_{\Delta})$, we obtain the following result.
Theorem 2.3.
Let $\varepsilon\in(0,1]$. Then, for every planar convex body $C\in\mathfrak{C}$ except for a meagre subset, there exist two infinite subsets $\mathcal{N}^{-}_{C},\mathcal{N}^{+}_{C}\subset\mathbb{N}$ such that
$\inf_{\#\mathcal{P}=N}\mathcal{D}_{2}(\mathcal{P},\,C)\leq\log^{1+\varepsilon}N\quad\forall N\in\mathcal{N}^{-}_{C}\quad\textnormal{ and }\quad\inf_{\#\mathcal{P}=N}\mathcal{D}_{2}(\mathcal{P},\,C)\geq N^{1/2-\varepsilon}\quad\forall N\in\mathcal{N}^{+}_{C}$.
3 The second method
The harmonic analysis arguments we are to employ in this section have been recently developed in [BT22] and [BER25]. These make use of a geometric quantity that we define as follows.
Figure 2: The chord in Definition 3.1.
Definition 3.1.
Let $C\subset\mathbb{R}^{2}$ be a convex body. For an angle $\theta\in[0,2\pi)$ and a value $\lambda>0$, we define the chord of $C$ in direction $\mathbf{u}(\theta)=(\cos\theta,\sin\theta)$ at distance $\lambda$ as
$K_{C}(\theta,\lambda)=\left\{\mathbf{x}\in C\,\colon\,\mathbf{x}\cdot\mathbf{u}(\theta)=\inf_{\mathbf{y}\in C}(\mathbf{y}\cdot\mathbf{u}(\theta))+\lambda\right\}.$
Further, we consider its length $\left|K_{C}(\theta,\lambda)\right|$, and we define the quantity
${\gamma}_{C}(\theta,\lambda)=\max\{\left|K_{C}(\theta,\lambda)\right|,\left|K_{C}(\theta+\pi,\lambda)\right|\}.$
One may notice that the decay of $|K_{C}(\theta,\lambda)|$ as $\lambda\to 0^{+}$ depends strongly on the local geometry of the (if just one) point $b\in\partial C$ whose inner normal is $\mathbf{u}(\theta)$. For example, if $\partial C$ is $\mathcal{C}^{2}$ and has positive curvature at $b$, then $|K_{C}(\theta,\lambda)|$ ultimately decays as $\lambda^{1/2}$, while if $b$ is the vertex at a corner, then $|K_{C}(\theta,\lambda)|$ ultimately decays as $\lambda$.
Now, we define the Fourier transform of $\mathds{1}_{C}$ as
$\widehat{\mathds{1}}_{C}(\boldsymbol{\xi})=\int_{C}e^{-2\pi i\mathbf{x}\cdot\boldsymbol{\xi}}\,\mathrm{d}\mathbf{x},$
and state a result that relates averages of the Fourier transform of $\mathds{1}_{C}$ to the chords of $C$.
Lemma 3.2.
Let $C$ be a planar convex body. Then, there exists a positive value $\lambda_{C}$ such that, for every angle $\theta\in[0,2\pi)$ and every $\lambda\in(0,\lambda_{C}]$, it holds
$\int_{0}^{1}\left|\widehat{\mathds{1}}_{\delta C}(\lambda^{-1}\,\mathbf{u}(\theta))\right|^{2}\,\mathrm{d}\delta\approx\lambda^{2}\gamma^{2}_{C}(\theta,\lambda).$
We refer to [BER25, Lem. 3.3] for a detailed proof of the latter, but we mention that it relies upon the one-dimensional results in [POD91] and [BT22].
We now show, through standard calculations, how quadratic discrepancy and the Fourier transform are intertwined. We start by exploiting the convolutional structure of (1.1). By setting
${\mu}=\sum_{\mathbf{p}\in\mathcal{P}}\mu_{\rm D}(-\mathbf{p})-N\mu_{\rm H},$
where $\mu_{\rm H}$ is the normalised Haar measure on $\mathbb{T}^{2}$ and $\mu_{\rm D}(\mathbf{p})$ is the Dirac delta centred at $\mathbf{p}$, one may write
$\mathcal{D}(\mathcal{P},\,\boldsymbol{\tau}+C)=\int_{\mathbb{T}^{2}}\mathfrak{P}\{\mathds{1}_{C}\}(-{\boldsymbol{\tau}}-\mathbf{x})\,\mathrm{d}{\mu}(\mathbf{x})=\left(\mathfrak{P}\{\mathds{1}_{C}\}\ast{\mu}\right)(-\boldsymbol{\tau}).$
Now, if $f$ is an integrable function or a finite measure on $\mathbb{T}^{2}$, we denote by
${\mathcal{F}}\{f\}\colon\mathbb{Z}^{2}\to\mathbb{C}$
the function of the Fourier coefficients of $f$. In particular, we note that it holds $\mathcal{F}\{{\mu}\}(\mathbf{0})=0$, and moreover, for every $\mathbf{n}\in\mathbb{Z}^{2}$, it holds
${\mathcal{F}}\circ\mathfrak{P}\{\mathds{1}_{C}\}(\mathbf{n})=\widehat{\mathds{1}}_{C}(\mathbf{n}).$
Hence, by applying Parseval’s identity, we obtain
$\int_{\mathbb{T}^{2}}\left|\mathcal{D}(\mathcal{P},\,\boldsymbol{\tau}+\delta C)\right|^{2}\,\mathrm{d}\boldsymbol{\tau}$ $=\int_{\mathbb{T}^{2}}\left|(\mathfrak{P}\{\mathds{1}_{\delta C}\}\ast{\mu})\right|^{2}(\boldsymbol{\tau})\,\mathrm{d}\boldsymbol{\tau}$
$=\sum_{{\mathbf{n}}\in\mathbb{Z}^{2}}\left|{\mathcal{F}}\circ\mathfrak{P}\{\mathds{1}_{\delta C}\}({\mathbf{n}})\right|^{2}\left|{\mathcal{F}}\{{\mu}\}({\mathbf{n}})\right|^{2}$
$=\sum_{{\mathbf{n}}\in\mathbb{Z}^{2}\setminus\{\mathbf{0}\}}\left|\widehat{\mathds{1}}_{\delta C}({\mathbf{n}})\right|^{2}\left|\sum_{\mathbf{p}\in\mathcal{P}}e^{2\pi i\mathbf{p}\cdot{\mathbf{n}}}\right|^{2}.$
Finally, we may rewrite (1.2) as
$\mathcal{D}_{2}(\mathcal{P},C)=\sum_{{\mathbf{n}}\in\mathbb{Z}^{2}\setminus\{\mathbf{0}\}}\left|\sum_{\mathbf{p}\in\mathcal{P}}e^{2\pi i\mathbf{p}\cdot{\mathbf{n}}}\right|^{2}\int_{0}^{1}\left|\widehat{\mathds{1}}_{\delta C}(\mathbf{n})\right|^{2}\,\mathrm{d}\delta.$ (3.1)
3.1 Geometric construction and Fourier estimates
Recalling the hypotheses of Theorem 1.6, we have that $\left\{\alpha_{i}\right\}_{i\in\mathbb{N}}$ is a sequence of exponents with values in $(2/5,1/2)$, $\left\{\varepsilon_{i}\right\}_{i\in\mathbb{N}}$ is a (possibly decreasing) sequence in $(0,\varepsilon]$ for some small positive $\varepsilon$, and $\left\{q_{i}\right\}_{i\in\mathbb{N}}$ is a (possibly increasing) sequence of positive integers. We remark that $\alpha_{i}$ prescribes the target order of growth, while $\varepsilon_{i}$ and $q_{i}$ quantify, respectively, how closely and for how long we approximate it.
We start our construction. Consider a planar convex body $C$ such that it has a centre of symmetry (this ensures $\gamma_{C}=|K_{C}|$) and is symmetric with respect to the $y$-axis. Then, suppose that $\mathbf{0}\in\partial C$ and that its inner normal is the vector $(0,1)$. By symmetry, we may construct $\partial C$ in the first quadrant $\mathbb{R}_{+}^{2}$. For this purpose, consider the graphs in $\mathbb{R}_{+}^{2}$ of the monomials
$\{P_{i}\}_{i\in\mathbb{N}}\quad\text{with}\quad P_{i}(x)=x^{\beta_{i}}\quad\text{and}\quad\beta_{i}=\frac{2\alpha_{i}}{2-3\alpha_{i}}\in(1,2),$
and notice that their curvatures
$\kappa_{i}(x)=\beta_{i}(\beta_{i}-1)\frac{x^{\beta_{i}-2}}{(1+\beta_{i}^{2}x^{2\beta_{i}-2})^{3/2}}$ (3.2)
are strictly decreasing on $(0,+\infty)$, and that $\kappa_{i}(x)\to+\infty$ as $x\to 0^{+}$.
Now, let $\left\{k_{i}\right\}_{i\in\mathbb{N}}$ be a sequence to be chosen later, but such that it is contained in $10,+\infty)$ and it holds $k_{i+1}>k_{i}$ with $k_{i}\to+\infty$. By induction, we construct a $\mathcal{C}^{2}$ curve segment of finite length as follows. First, consider the auxiliary curve segment
$\left\{(x,y)\in\mathbb{R}_{+}^{2}\,\colon\,P_{1}(x)=y,\,a_{1}<x<b_{1},\,\kappa_{1}(a_{1})=k_{1},\,\kappa_{1}(b_{1})=10\right\},$
and then glue at the endpoint $(a_{1},P_{1}(a_{1}))$ a tangentially-aligned copy of the auxiliary curve segment
$\left\{(x,y)\in\mathbb{R}_{+}^{2}\,\colon\,P_{2}(x)=y,\,a_{2}<x<b_{2},\,\kappa_{2}(a_{2})=k_{2},\,\kappa_{2}(b_{2})=k_{1}\right\}.$
Figure 3: A close-up look at $P$, in which different auxiliary monomial curve segments are displayed in different colours.
In particular, notice that we obtain a $\mathcal{C}^{2}$ link since the signed curvatures at the endpoints $(a_{1},P_{1}(a_{1}))$ and $(b_{2},P_{2}(b_{2}))$ coincide. Inductively, and in such a $\mathcal{C}^{2}$ fashion, we glue all the auxiliary curve segments
$\left\{(x,y)\in\mathbb{R}_{+}^{2}\,\colon\,P_{i}(x)=y,\,a_{i}<x<b_{i},\,\kappa_{i}(a_{i})=k_{i},\,\kappa_{i}(b_{i})=k_{i-1}\right\},$
and in the end, we obtain a $\mathcal{C}^{2}$ curve segment of finite length. By a final rigid motion, we may place this curve segment so that it is tangent to the $x$-axis and its left endpoint is the origin; see Figure 3 for an illustration of the resulting curve segment $P$. By construction, $P$ has strictly decreasing continuous curvature (and this goes to infinity as one approaches the origin). Finally, we let $\partial C$ coincide with $P$ for its length, and then we design the rest of $\partial C$ in such a way that it is $\mathcal{C}^{2}$ with signed curvature bounded from above by $\kappa_{1}(1)$ and from below by, say, $1/10$.
The purpose of the latter construction is the following: $\partial C$ has to look like the graphs of different $P_{i}$’s at increasingly smaller neighbourhoods of the origin. Ultimately, this is achieved by ensuring that the sequence $\left\{k_{i}\right\}_{i\in\mathbb{N}}$ increases very rapidly. Indeed, for every $i\in\mathbb{N}$, mark by $(x_{i},y_{i})$ the left endpoint of $P_{i}$ after being assembled into $P$, as in Figure 3. First, the fact that, for $i\in\mathbb{N}$ and $x\in(0,1]$, it holds
$\kappa_{i}(x)\leq 2x^{-1}\quad\text{and}\quad P^{\prime}_{i}(x)\leq 2,$
ensures that we may get any $(x_{i},y_{i})$ arbitrarily close to the origin by a large enough choice of $k_{i}$. Secondly, since the elements in $\left\{k_{j}\right\}_{j\geq i}$ may be chosen arbitrarily large, we may get the elements in $\left\{P_{j+1}^{\prime}(b_{j+1})\right\}_{j\geq i}$ arbitrarily close to 0, and this, in turn, ensures that we may get the inner normal at $(x_{i},y_{i})$ arbitrarily close to the vector $(0,1)$; in particular, this prevents the curve segment $P$ from becoming a spiral.
In order to apply Lemma 3.2 to $C$, we are interested in how the chords $K_{C}(\theta+\pi/2,\lambda)$ behave for small values of $\theta$. For this purpose, let $\beta\in(1,2)$, and consider a second planar convex body $A(\beta)$ whose boundary coincides with the graph of $y=|x|^{\beta}$ in $1,1]^{2}$. Moreover, suppose that $A(\beta)$ has a centre of symmetry and is symmetric with respect to the $y$-axis. Hence, it holds the following estimates on the chords of $A(\beta)$.
Proposition 3.3.
Let $A(\beta)$ be as previously defined. Then, there exists a positive small constant $c$ (independent of $\beta$) such that it holds
$\left|K_{A(\beta)}(\theta+\pi/2,\lambda)\right|\approx_{\beta}\begin{cases}\lambda^{1/\beta}&\textnormal{if}\quad 0\leq|\theta|<\lambda^{\frac{\beta-1}{\beta}}<c\\ \lambda^{1/2}|\theta|^{\frac{2-\beta}{2(\beta-1)}}&\textnormal{if}\quad\lambda^{\frac{\beta-1}{\beta}}\leq|\theta|<c\end{cases}.$
We leave the proof of the latter in the Appendix, and we mention that similar calculations had previously been carried out in [BER25, Prop. 8.4].
Recall that, by a large enough choice of the elements in $\left\{k_{j}\right\}_{j\geq i}$, we may get $(x_{i},y_{i})$ arbitrarily close to the origin, as we may get the inner normal at $(x_{i},y_{i})$ to be arbitrarily close to the vector $(0,1)$. Ultimately, this guarantees that, for every $i\in\mathbb{N}$, we may obtain an interval $\lambda_{i,1},\lambda_{i,2}]$ and a positive value $\theta_{i}$ such that, for every $\theta$ such that $|\theta|\leq\theta_{i}$ and every $\lambda$ such that $\lambda\in\left[\lambda_{i,1},\lambda_{i,2}\right]$, it holds
$\left|K_{C}(\theta+\pi/2,\lambda)\right|\approx\left|K_{A(\beta_{i})}(\theta+\pi/2,\lambda)\right|.$
Notice that, as $\theta_{i}$ and $\lambda_{i,2}$ depend on $\left\{\beta_{j}\right\}_{j=1}^{i-1}$ and $\left\{k_{j}\right\}_{j=1}^{i-1}$, the choice of $\lambda_{i,1}$ can be made arbitrarily small at every $i$-th step; this will be key for the estimates in the following subsection.
Finally, Proposition 3.3 and Lemma 3.2 may be applied to $C$ to infer the following result. From now on, we write $\rho^{-1}$ in place of $\lambda$, to prioritise the argument of the Fourier transform.
Lemma 3.4.
Let $C$ be as previously defined. Then, one may choose the sequence $\left\{k_{i}\right\}_{i\in\mathbb{N}}$ in such a way that, for every $i\in\mathbb{N}$, there exists an interval $\rho_{i,1},\rho_{i,2}]\subset[1,+\infty)$ (with $\rho_{i+1,1}>\rho_{i,2}$ and $\lim_{i\to+\infty}\rho_{i,1}=+\infty$) and a positive value $\theta_{i}$ (with $\theta_{i+1}<\theta_{i}$ and $\lim_{i\to+\infty}\theta_{i}=0$) such that, for $\rho\in[\rho_{i,1},\rho_{i,2}]$ and $|\theta|\leq\theta_{i}$, it holds
$\int_{0}^{1}\left|\widehat{\mathds{1}}_{\delta C}\left(\rho\,\mathbf{u}(\theta+\pi/2)\right)\right|^{2}\,\mathrm{d}\delta\approx_{\beta_{i}}\begin{cases}\rho^{-2-2/\beta_{i}}&\textnormal{if}\quad\left|\theta\right|\leq\rho^{\frac{1-\beta_{i}}{\beta_{i}}}\\ \rho^{-3}\left|\theta\right|^{\frac{2-\beta_{i}}{\beta_{i}-1}}&\textnormal{if}\quad\rho^{\frac{1-\beta_{i}}{\beta_{i}}}<\left|\theta\right|\leq\theta_{i}\end{cases}.$
In particular, although $\rho_{i,1}$ and $\theta_{i}$ are determined by the elements in $\left\{k_{j}\right\}_{j=1}^{i-1}$, the endpoint $\rho_{i,2}$ can be made arbitrarily large upon a larger choice of $k_{i}$.
Starting from the fact that we have designed $P$ so that its curvature is strictly decreasing, we obtain the following auxiliary result.
Lemma 3.5.
Let $C$ be as previously defined. Then, there exist $\lambda_{0}>0$ and $\theta_{0}>\pi/2$ such that, for $\lambda\in[0,\lambda_{0})$ and for $\pi/2\leq\theta_{1}<\theta_{2}<\theta_{0}$, it holds
$|K_{C}(\theta_{1},\lambda)|\leq 2|K_{C}(\theta_{2},\lambda)|.$
The proof of the latter requires a technical geometric argument, which we leave to the Appendix. By pairing Lemma 3.2 with Lemma 3.5, we obtain that, for every $\rho>\max\{\lambda_{C}^{-1},\lambda_{0}^{-1}\}$ and every angle $\theta_{1}$ and $\theta_{2}$ such that $\pi/2\leq\theta_{1}<\theta_{2}<\theta_{0}$, it holds
$\int_{0}^{1}\left|\widehat{\mathds{1}}_{\delta C}\left(\rho\,\mathbf{u}(\theta_{1})\right)\right|^{2}\,\mathrm{d}\delta\lesssim\int_{0}^{1}\left|\widehat{\mathds{1}}_{\delta C}\left(\rho\,\mathbf{u}(\theta_{2})\right)\right|^{2}\,\mathrm{d}\delta.$ (3.3)
3.2 Lower bound
We show how we may construct the sequence $\left\{k_{i}\right\}_{i\in\mathbb{N}}$ in such a way that, for every $i\in\mathbb{N}$, it holds the lower bound
$N^{\alpha_{i}-\varepsilon_{i}}\leq\inf_{\#\mathcal{P}=N}\mathcal{D}_{2}(\mathcal{P},\,C)\quad\text{for}\quad N\in[N_{i},N_{i}+q_{i}].$
For this purpose, we need an argument of Cassels [CAS56] and Montgomery [MON94, Ch. 6] for estimating exponential sums from below. We formulate it in a version specialised to balls, and the reader may find a short proof either in [BT22] or [BER25]. In what follows, $B(r)$ stands for the ball centred at the origin with radius $r$.
Lemma 3.6 (Cassels-Montgomery).
For every origin-symmetric convex body $C\subset\mathbb{R}^{2}$ and for every finite configuration of points $\{\mathbf{p}_{j}\}_{j=1}^{N}$ in $\mathbb{T}^{2}$, it holds
$\sum_{\mathbf{m}\in(C\setminus B(r))\cap\mathbb{Z}^{2}}\left|\sum_{j=1}^{N}e^{2\pi i\mathbf{m}\cdot\mathbf{p}_{j}}\right|^{2}\geq\frac{|C|}{4}N-2\pi(r^{2}+1)N^{2}.$
We proceed with the geometric argument that will pair with the latter lemma. For the sake of notation, we write
$\|\theta\|_{\pi}=\min_{j\in\mathbb{Z}}|\theta-j\pi|.$
First, notice that, since we have constructed $\partial C$ in such a way that its curvature is uniformly bounded from below and above when excluding a neighbourhood of the origin (and its symmetric counterpart), then for every $\lambda\in[0,\lambda_{0})$ and for every $\theta$ such that $\|\theta\|_{\pi}>1/10$, it holds
$|K_{C}(\theta+\pi/2,\lambda)|\approx\lambda^{1/2}.$ (3.4)
Now, fix an $i\in\mathbb{N}$. Then, by the latter consideration paired with Lemma 3.2, and by Lemma 3.4 and (3.3), we get the lower bound
$\int_{0}^{1}\left|\widehat{\mathds{1}}_{\delta C}\left(\rho\,\mathbf{u}(\theta+\pi/2)\right)\right|^{2}\,\mathrm{d}\delta\gtrsim_{\beta_{i}}\begin{cases}\rho^{-2-2/\beta_{i}}&\textnormal{if}\quad\rho\in\left[\rho_{i,1},\rho_{i,2}\right]\;\land\;\|\theta\|_{\pi}\leq 1/10\\ \rho^{-3}&\textnormal{if}\quad\rho\in\left[\rho_{i,1},\rho_{i,2}\right]\;\land\;\|\theta\|_{\pi}>1/10\\ 0&\textnormal{else}\end{cases}.$
Let $\tilde{\varepsilon}_{i}\in(0,1]$ be a value to be chosen later. Moreover, let $N$, $X$, and $Y$ be positive parameters to be chosen later, but such that $\rho_{i,1}\ll Y\ll X\ll\rho_{i,2}$ and $XY=N^{1+\tilde{\varepsilon}_{i}}$. Then, consider a rectangle $R\subset\mathbb{R}^{2}$ such that it is symmetric with respect to the axes and has a vertex in $(X/2,Y/2)$; consequently, it also holds $|R|=N^{1+\tilde{\varepsilon}_{i}}$. Denote by $R(\omega)$ the same rectangle rotated counterclockwise by an angle $\omega$, and define the function $\Phi\colon\mathbb{Z}^{2}\to[0,+\infty)$ by
$\Phi(\mathbf{m})=\int_{-1/10}^{1/10}\mathds{1}_{R(\omega)\setminus B(\rho_{i,1})}(\mathbf{m})\,\mathrm{d}\omega,$
Now, in order to apply the Cassels-Montgomery lemma effectively, we need to find a positive parameter $Z=Z(N)$ such that in the annulus $|\mathbf{m}|\in\left[\rho_{i,1},\rho_{i,2}\right]$ it holds
$Z\Phi(\mathbf{m})\leq\begin{cases}|\mathbf{m}|^{-2-2/\beta_{i}}&\textnormal{if}\quad|\mathbf{m}|\in\left[\rho_{i,1},\rho_{i,2}\right]\;\land\;\|\arg(\mathbf{m})+\pi/2\|_{\pi}\leq 1/10\\ |\mathbf{m}|^{-3}&\textnormal{if}\quad|\mathbf{m}|\in\left[\rho_{i,1},\rho_{i,2}\right]\;\land\;\|\arg(\mathbf{m})+\pi/2\|_{\pi}>1/10\\ 0&\textnormal{else}\end{cases}.$
First, we describe two regions where $\Phi$ vanishes. By construction, for all $\mathbf{m}\in\mathbb{Z}^{2}$ such that $|\mathbf{m}|\in\left[\rho_{i,1},X\right]^{\mathsf{c}}$, it holds $\Phi(\mathbf{m})=0$. Secondly, for all $\mathbf{m}\in\mathbb{Z}^{2}$ such that $|\mathbf{m}|\geq Y$ and $\|\arg(\mathbf{m})+\pi/2\|_{\pi}\leq 1/10$, it holds $\Phi(\mathbf{m})=0$ as well.
Now, for all $\mathbf{m}\in\mathbb{Z}^{2}$ such that $|\mathbf{m}|\leq Y$, we have the trivial estimate $\Phi(\mathbf{m})\leq 1/5$. Since in this region we are aiming for $Z\Phi(\mathbf{m})\leq|\mathbf{m}|^{-2-2/\beta_{i}}$, then we settle for
$Z\lesssim Y^{-2-2/\beta_{i}}.$
Lastly, we consider the remaining region defined by $|\mathbf{m}|\in[Y,X]$ and $\|\arg(\mathbf{m})+\pi/2\|_{\pi}>1/10$. By some elementary geometry, we find that $\Phi(\mathbf{m})\leq 2\psi$ where $\psi$ is such that $|\mathbf{m}|\sin\psi=Y/2$. Hence, it holds $\Phi(\mathbf{m})\leq\pi Y/(2|\mathbf{m}|)$, and since in this region we are aiming for $Z\Phi(\mathbf{m})\leq|\mathbf{m}|^{-3}$, then we settle for
$Z\lesssim Y^{-1}X^{-2}.$
Having considered all the possible regions, we are left with the requirement
$Z\lesssim\min\left(Y^{-2-2/\beta_{i}},Y^{-1}X^{-2}\right),$
and if, by an optimisation argument, we equalise the two terms in the minimum and use the constraint $XY=N^{1+\tilde{\varepsilon}_{i}}$, then we obtain
$X\approx N^{(1+\tilde{\varepsilon}_{i})(\beta_{i}+2)/(3\beta_{i}+2)},\quad Y\approx N^{(1+\tilde{\varepsilon}_{i})2\beta_{i}/(3\beta_{i}+2)},\quad\text{and}\quad Z\approx N^{-(1+\tilde{\varepsilon}_{i})(4\beta_{i}+4)/(3\beta_{i}+2)}.$
Starting from (3.1), and by Parseval’s identity and the Cassels-Montgomery lemma, we get that, for any configuration of $N$ points $\mathcal{P}=\{\mathbf{p}_{j}\}_{j=1}^{N}$ in $\mathbb{T}^{2}$, it holds
$\int_{0}^{1}\int_{\mathbb{T}^{2}}\left|\mathcal{D}(\mathcal{P},\,\boldsymbol{\tau}+\delta C)\right|^{2}\,\mathrm{d}\boldsymbol{\tau}\,\mathrm{d}\delta$ $=\sum_{\mathbf{m}\in\mathbb{Z}^{2}\setminus\{\mathbf{0}\}}\left|\sum_{j=1}^{N}e^{2\pi i\mathbf{m}\cdot\mathbf{p}_{j}}\right|^{2}\int_{0}^{1}\left|\widehat{\mathds{1}}_{\delta C}\left(\mathbf{m}\right)\right|^{2}\,\mathrm{d}\delta$
$\gtrsim_{\beta_{i}}\sum_{|\mathbf{m}|\geq\rho_{i,1}}\left|\sum_{j=1}^{N}e^{2\pi i\mathbf{m}\cdot\mathbf{p}_{j}}\right|^{2}Z\Phi(\mathbf{m})$
$=Z\int_{-1/10}^{1/10}\sum_{\mathbf{m}\in R(\omega)\setminus B(\rho_{i,1})}\left|\sum_{j=1}^{N}e^{2\pi i\mathbf{m}\cdot\mathbf{p}_{j}}\right|^{2}\,\mathrm{d}\omega$
$\gtrsim Z\left(N^{2+\tilde{\varepsilon}_{i}}-2\pi(\rho_{i,1}^{2}+1)N^{2}\right).$
By construction, $\rho_{i,2}$ can be chosen arbitrarily large, and this also applies to $N$ since our only constraint is $X\ll\rho_{i,2}$. Hence, by taking $\tilde{\varepsilon}_{i}$ in such a way that $\tilde{\varepsilon}_{i}(\beta_{i}+2)/(3\beta_{i}+2)<\varepsilon_{i}$, we find an $N_{i}$ such that
$N_{i}^{\tilde{\varepsilon}_{i}}\geq 4\pi(\rho^{2}_{i,1}+1)\quad\text{and}\quad(N_{i}+q_{i})^{(1+\tilde{\varepsilon}_{i})(\beta_{i}+2)/(3\beta_{i}+2)}\ll\rho_{i,2}.$
Therefore, it follows that, for every $N\in[N_{i},N_{i}+q_{i}]$, it holds
$\int_{0}^{1}\int_{\mathbb{T}^{2}}\left|\mathcal{D}(\mathcal{P},\,\boldsymbol{\tau}+\delta C)\right|^{2}\,\mathrm{d}\boldsymbol{\tau}\,\mathrm{d}\delta\gtrsim_{\beta_{i}}N^{-(1+\tilde{\varepsilon}_{i})(4\beta_{i}+4)/(3\beta_{i}+2)}N^{2+\tilde{\varepsilon}_{i}}=N^{\frac{2\beta_{i}}{3\beta_{i}+2}-\tilde{\varepsilon}_{i}\frac{\beta_{i}+2}{3\beta_{i}+2}}>N^{\alpha_{i}-\varepsilon_{i}}.$
Notice that, since $N^{-\varepsilon_{i}}$ can be made arbitrarily small, we can ultimately get the implicit constant in front of $N^{\alpha_{i}-\varepsilon_{i}}$ to be $1$.
3.3 Upper bound
For every $i\in\mathbb{N}$, we proceed to show the upper bound
$\inf_{\#\mathcal{P}=N}\mathcal{D}_{2}(\mathcal{P},\,C)\lesssim N^{\alpha_{i}+\varepsilon_{i}}\quad\text{for}\quad N\in[N_{i},N_{i}+q_{i}].$
This is done by employing suitable sampling sequences, and we will make use of the kind that appear in [BT22, Thm. 8].
Fix $i\in\mathbb{N}$, and again, we recall that at every $i$-th step of our construction, $\rho_{i,2}$ serves as a free parameter that may be chosen arbitrarily large. By pairing Lemma 3.2 with (3.4), and by Lemma 3.4 and (3.3), in the annulus $\rho\in[\rho_{i,1},\rho_{i,2}]$ it holds the upper bound
$\int_{0}^{1}\left|\widehat{\mathds{1}}_{\delta C}\left(\rho\,\mathbf{u}(\theta+\pi/2)\right)\right|^{2}\,\mathrm{d}\delta\lesssim_{\beta_{i}}\begin{cases}\rho^{-2-2/\beta_{i}}&\textnormal{if}\quad\|\theta\|_{\pi}\leq\rho^{\frac{1-\beta_{i}}{\beta_{i}}}\\ \rho^{-3}\|\theta\|_{\pi}^{\frac{2-\beta_{i}}{\beta_{i}-1}}&\textnormal{if}\quad\rho^{\frac{1-\beta_{i}}{\beta_{i}}}<\|\theta\|_{\pi}\leq\theta_{i}\\ \rho^{-3}&\textnormal{if}\quad\theta_{i}<\|\theta\|_{\pi}\par\end{cases}.$ (3.5)
By Lemma 3.2, and since we have constructed $\partial C$ in such a way that its curvature is uniformly bounded from below, then, for every $\rho>\lambda_{C}^{-1}$ and any angle $\theta$, it holds
$\int_{0}^{1}\left|\widehat{\mathds{1}}_{\delta C}\left(\rho\,\mathbf{u}(\theta)\right)\right|^{2}\,\mathrm{d}\delta\lesssim\rho^{-3}.$ (3.6)
Again, let $\tilde{N}\in\mathbb{N}$ be a free large parameter to be chosen later, and set
$G=\lfloor\tilde{N}^{\frac{2+\beta_{i}}{2+3\beta_{i}}}\rfloor,\quad L=\lfloor\tilde{N}^{\frac{2\beta_{i}}{2+3\beta_{i}}}\rfloor,\quad J_{G}=[0,G-1]\cap\mathbb{N},\quad\text{and}\quad J_{L}=[0,L-1]\cap\mathbb{N}.$
Set $N=GL$ (trivially, $L<G$ and $N/2<\tilde{N}<2N$), and consider the configuration of $N$ points $\mathcal{P}$ in $\mathbb{T}^{2}$ defined by
$\mathcal{P}=\left\{\mathbf{p}_{j}\right\}_{j=1}^{N}=\{\mathbf{p}_{\ell,g}\}_{g\in J_{G},\,\ell\in J_{L}}\quad\text{with}\quad\mathbf{p}_{\ell,g}=\left(\frac{g}{G},\,\frac{\ell}{L}\right).$
Recall that, by Parseval’s identity, we get
$\int_{\mathbb{T}^{2}}\left|\mathcal{D}(\mathcal{P},\boldsymbol{\tau}+\delta C)\right|^{2}\,\mathrm{d}\boldsymbol{\tau}=\sum_{\mathbf{m}\neq(0,0)}\left|\widehat{\mathds{1}}_{\delta C}(\mathbf{m})\right|^{2}\left|\sum_{g\in J_{G}}\sum_{\ell\in J_{L}}e^{2\pi i\mathbf{m}\cdot\mathbf{p}_{\ell,g}}\right|^{2},$
and in particular, it holds the identity
$\sum_{g\in J_{G}}\sum_{\ell\in J_{L}}e^{2\pi i\left(m_{1}\frac{g}{G}+m_{2}\frac{\ell}{L}\right)}=\begin{cases}GL&\text{if}\quad m_{1}\in G\mathbb{Z}\quad\text{and}\quad m_{2}\in L\mathbb{Z}\\ 0&\text{else}\end{cases}.$
Hence, consider the set $\mathcal{R}=(G\mathbb{Z}\times L\mathbb{Z})\setminus\{\mathbf{0}\}$, and split it into the regions
$\begin{split}V_{1}&=\left\{\mathbf{m}\in\mathcal{R}\,\colon\,|m_{1}|\leq|m_{2}|^{1/\beta_{i}}\right\},\\ V_{2}&=\left\{\mathbf{m}\in\mathcal{R}\,\colon\,|m_{2}|^{1/\beta_{i}}<|m_{1}|\leq\theta_{i}|m_{2}|\right\},\\ V_{3}&=\left\{\mathbf{m}\in\mathcal{R}\,\colon\,\theta_{i}|m_{2}|<|m_{1}|\right\},\end{split}$
so that we may write
$\mathcal{D}_{2}(\mathcal{P},\,C)=G^{2}L^{2}\left(\sum_{\mathbf{m}\in V_{1}}+\sum_{\mathbf{m}\in V_{2}}+\sum_{\mathbf{m}\in V_{3}}\right)\int_{0}^{1}\left|\widehat{\mathds{1}}_{\delta C}(\mathbf{m})\right|^{2}\,\mathrm{d}\delta.$ (3.7)
Now, we constrain $\tilde{N}$ to be so large that $L>\rho_{i,1}$, and as a consequence, it follows that $\mathcal{R}\cap B(\rho_{i,1})=\varnothing$. Moreover, we constrain $\rho_{i,2}$ to be so large that
$\mathcal{Q}_{i}=\sum_{\mathbf{m}\in\mathcal{R}\cap B^{\mathsf{c}}(\rho_{i,2})}|\mathbf{m}|^{-3}<N^{-2}.$
We exploit (3.5) to study the three sums in (3.7). In this case, we must consider
$\rho=|\mathbf{m}|\quad\text{and}\quad\tan\theta=-\frac{m_{1}}{m_{2}}.$
For the sake of notation, we write the annulus $\mathcal{A}=\left\{\mathbf{m}\in\mathcal{R}\,\colon\,|\mathbf{m}|\in[\rho_{i,1},\rho_{i,2}]\right\}$. Notice that for small $\theta$ it holds $\tan\theta\approx\theta$, and consequently, with a bit of rearrangement, we can rewrite the estimates in (3.5) as: for every $\mathbf{m}\in\mathcal{A}$ it holds
$\int_{0}^{1}\left|\widehat{\mathds{1}}_{\delta C}\left(\mathbf{m}\right)\right|^{2}\,\mathrm{d}\delta\lesssim_{\beta_{i}}\begin{cases}|m_{2}|^{-2-2/\beta_{i}}&\text{if}\quad|m_{1}|\leq|m_{2}|^{1/\beta_{i}}\\ |m_{1}|^{\frac{2-\beta_{i}}{\beta_{i}-1}}|m_{2}|^{\frac{1-2\beta_{i}}{\beta_{i}-1}}&\text{if}\quad|m_{2}|^{1/\beta_{i}}\leq|m_{1}|<\theta_{i}|m_{2}|\\ |m_{1}|^{-3}&\text{if}\quad\theta_{i}|m_{2}|<|m_{1}|\end{cases}.$ (3.8)
For $j\in\{1,2,3\}$, notice that by the constraints we have imposed earlier, it holds $V_{j}\cap\mathcal{A}=V_{j}\cap B(\rho_{i,2})$ and $G^{2}L^{2}\mathcal{Q}_{i}=N^{2}\mathcal{Q}_{i}<1$, and by further applying (3.6), we obtain
$\begin{split}G^{2}L^{2}\sum_{\mathbf{m}\in V_{j}}\int_{0}^{1}\left|\widehat{\mathds{1}}_{\delta C}(\mathbf{m})\right|^{2}\,\mathrm{d}\delta&=G^{2}L^{2}\left(\sum_{\mathbf{m}\in V_{j}\cap\mathcal{A}}+\sum_{\mathbf{m}\in V_{j}\cap B^{\mathsf{c}}(\rho_{i,2})}\right)\int_{0}^{1}\left|\widehat{\mathds{1}}_{\delta C}(\mathbf{m})\right|^{2}\,\mathrm{d}\delta\\ &\lesssim G^{2}L^{2}\sum_{\mathbf{m}\in V_{j}\cap\mathcal{A}}\int_{0}^{1}\left|\widehat{\mathds{1}}_{\delta C}(\mathbf{m})\right|^{2}\,\mathrm{d}\delta+G^{2}L^{2}\mathcal{Q}_{i}\\ &\lesssim G^{2}L^{2}\sum_{\mathbf{m}\in V_{j}\cap\mathcal{A}}\int_{0}^{1}\left|\widehat{\mathds{1}}_{\delta C}(\mathbf{m})\right|^{2}\,\mathrm{d}\delta.\end{split}$ (3.9)
Then, by applying (3.8) and (3.9), for the first sum in the last term of (3.7) we obtain
$\sum_{\mathbf{m}\in V_{1}}\int_{0}^{1}\left|\widehat{\mathds{1}}_{\delta C}(\mathbf{m})\right|^{2}\,\mathrm{d}\delta$ $\lesssim_{\beta_{i}}\sum_{\mathbf{m}\in V_{1}}|m_{2}|^{-2-2/\beta_{i}}\lesssim\sum_{n_{2}=1}^{+\infty}\,\sum_{n_{1}=0}^{(Ln_{2})^{1/\beta_{i}}/G}(Ln_{2})^{-2-2/\beta_{i}}$
$\lesssim L^{-2-2/\beta_{i}}\sum_{n_{2}=1}^{+\infty}n_{2}^{-2-2/\beta_{i}}\left(1+(Ln_{2})^{1/\beta_{i}}/G\right)$
$\lesssim L^{-2-2/\beta_{i}}.$
Similarly, for the second sum in the last term of (3.7), we get
$\sum_{\mathbf{m}\in V_{2}}\int_{0}^{1}\left|\widehat{\mathds{1}}_{\delta C}(\mathbf{m})\right|^{2}\,\mathrm{d}\delta$ $\lesssim_{\beta_{i}}\sum_{\mathbf{m}\in V_{2}}|m_{1}|^{\frac{2-\beta_{i}}{\beta_{i}-1}}|m_{2}|^{\frac{1-2\beta_{i}}{\beta_{i}-1}}$
$\lesssim\sum_{n_{1}=1}^{+\infty}\,\sum_{n_{2}=Gn_{1}/(\theta_{i}L)}^{(Gn_{1})^{\beta_{i}}/L}(Gn_{1})^{\frac{2-\beta_{i}}{\beta_{i}-1}}(Ln_{2})^{\frac{1-2\beta_{i}}{\beta_{i}-1}}$
$\lesssim G^{\frac{2-\beta_{i}}{\beta_{i}-1}}L^{\frac{1-2\beta_{i}}{\beta_{i}-1}}\sum_{n_{1}=1}^{+\infty}n_{1}^{\frac{2-\beta_{i}}{\beta_{i}-1}}\sum_{n_{2}=Gn_{1}/(\theta_{i}L)}^{+\infty}n_{2}^{\frac{1-2\beta_{i}}{\beta_{i}-1}}$
$\lesssim G^{\frac{2-\beta_{i}}{\beta_{i}-1}}L^{\frac{1-2\beta_{i}}{\beta_{i}-1}}\sum_{n_{1}=1}^{+\infty}n_{1}^{\frac{2-\beta_{i}}{\beta_{i}-1}}\left(Gn_{1}/(\theta_{i}L)\right)^{\frac{-\beta_{i}}{\beta_{i}-1}}$
$\lesssim G^{-2}L^{-1}\sum_{n_{1}=1}^{+\infty}n_{1}^{-2}\lesssim G^{-2}L^{-1}.$
For the last sum in the last term in (3.7), we obtain
$\sum_{\mathbf{m}\in V_{3}}\int_{0}^{1}\left|\widehat{\mathds{1}}_{\delta C}(\mathbf{m})\right|^{2}\,\mathrm{d}\delta$ $\lesssim_{\beta_{i}}\sum_{\mathbf{m}\in V_{3}}|m_{1}|^{-3}\lesssim\sum_{n_{1}=1}^{+\infty}\,\sum_{n_{2}=0}^{Gn_{1}/(\theta_{i}L)}(Gn_{1})^{-3}$
$\lesssim G^{-3}\sum_{n_{1}=1}^{+\infty}n_{1}^{-2}G/(\theta_{i}L)\lesssim G^{-2}L^{-1}\theta_{i}^{-1}.$
Finally, since $G^{2}L^{-2/\beta_{i}}=L\approx N^{\frac{2\beta_{i}}{2+3\beta_{i}}}=N^{\alpha_{i}}$, and by comparing the previous three estimates with (3.7), we obtain
$\mathcal{D}_{2}(\mathcal{P},\,C)\lesssim_{\beta_{i}}G^{2}L^{2}\left(L^{-2-2/\beta_{i}}+G^{-2}L^{-1}/\theta_{i}\right)\lesssim G^{2}L^{-2/\beta_{i}}+L/\theta_{i}\lesssim N^{\alpha_{i}}\theta_{i}^{-1}$
Hence, since we may choose $\rho_{i,2}$ and $\tilde{N}$ arbitrarily large, then we can ultimately obtain $N$ such that $N^{\varepsilon_{i}}\geq\theta_{i}^{-1}$, and for the same reason, we may ultimately write
$\mathcal{D}_{2}(\mathcal{P},\,C)\leq N^{\alpha_{i}+\varepsilon_{i}}.$
At this stage, by the constraints we have imposed, we may only conclude that the upper bound holds for all the integers $N$ of the form
$N=\lfloor n^{\frac{2+\beta_{i}}{2+3\beta_{i}}}\rfloor\,\lfloor n^{\frac{2\beta_{i}}{2+3\beta_{i}}}\rfloor,\quad\text{with}\quad n^{\alpha_{i}}>\rho_{i,1}\quad\text{and}\quad n^{2}<\rho_{i,2}.$
In order to prove that there is a suitable choice of points for some $q_{i}$ consecutive integers in the interval $(\rho_{i,1}^{1/\alpha_{i}},\rho_{i,2}^{1/2})$, consider the following recursive definition
$n_{j}=\max\left\{n\in\mathbb{N}\;\colon\lfloor n^{1-\alpha_{i}}\rfloor\,\lfloor n^{\alpha_{i}}\rfloor\leq N-\sum_{k=1}^{j-1}\,\lfloor n_{k}^{1-\alpha_{i}}\rfloor\,\lfloor n_{k}^{\alpha_{i}}\rfloor\right\}\quad\text{for}\quad j\in\mathbb{N}\setminus\{0\},$
where improper sums are conventionally considered as zeros. Now, notice that the latter definition implies that
$\begin{split}N-\sum_{k=1}^{j-1}\,\lfloor n_{k}^{1-\alpha_{i}}\rfloor\,\lfloor n_{k}^{\alpha_{i}}\rfloor&\leq\lfloor(n_{j}+1)^{1-\alpha_{i}}\rfloor\,\lfloor(n_{j}+1)^{\alpha_{i}}\rfloor\\ &\leq\lfloor n_{j}^{1-\alpha_{i}}\rfloor\,\lfloor n_{j}^{\alpha_{i}}\rfloor+\lfloor n_{j}^{1-\alpha_{i}}\rfloor+\lfloor n_{j}^{\alpha_{i}}\rfloor+1,\end{split}$
and therefore, it follows that
$N-\sum_{k=1}^{j}\,\lfloor n_{k}^{1-\alpha_{i}}\rfloor\,\lfloor n_{k}^{\alpha_{i}}\rfloor\leq 2n_{j}^{1-\alpha_{i}}.$
Again, by the definition of $n_{j}$, we point out that
$\frac{n_{1}}{2}\leq N\quad\text{and}\quad\frac{n_{j+1}}{2}\leq N-\sum_{k=1}^{j}\,\lfloor n_{k}^{1-\alpha_{i}}\rfloor\,\lfloor n_{k}^{\alpha_{i}}\rfloor,$
and therefore, by induction, we obtain
$N-\sum_{k=1}^{j}\,\lfloor n_{k}^{1-\alpha_{i}}\rfloor\,\lfloor n_{k}^{\alpha_{i}}\rfloor\leq 2n_{j}^{1-\alpha_{i}}\leq 2^{4}n_{j-1}^{(1-\alpha_{i})^{2}}\leq 2^{2j}N^{\left(1-\alpha_{i}\right)^{j}}.$
In particular, notice that
$N-\sum_{k=1}^{5}\lfloor n_{k}^{1-\alpha_{i}}\rfloor\,\lfloor n_{k}^{\alpha_{i}}\rfloor\leq 2^{10}N^{\left(1-\alpha_{i}\right)^{5}}\ll\bigl(N^{\alpha_{i}/2}\bigr).$
Then, if we further impose that $N^{\alpha_{i}^{2}/2}>\rho_{i,1}$, we may associate a choice of points as in the previous construction to every $N_{j}=\lfloor n_{j}^{1-\alpha_{i}}\rfloor\,\lfloor n_{j}^{\alpha_{i}}\rfloor$ such that $N_{j}\gtrsim N^{\alpha_{i}/2}$, and we do not bother the contribution of the remaining points since, by the latter equation, it is certainly much smaller than $N^{\alpha_{i}}$. In particular, it is not restrictive to admit repetition of points in the sampling sequence $\mathcal{P}_{N}$. Finally, it is harmless to sum the contribution of our (at most) five point-configurations and the remainder points, since, for $a_{1},\ldots,a_{6}\geq 0$, it holds
$\left(\sum_{j=1}^{6}a_{j}\right)^{2}\leq 6\sum_{j=1}^{6}a_{j}^{2}.$
By the argument we have just carried over, we may find an integer $N_{i}\gg q_{i}$ for which the upper bound in Theorem 1.6 holds, and by choosing $\rho_{i,2}$ large enough, we ensure that the same construction may be carried over for the next consecutive $q_{i}$ integers after $N_{i}$.
Remark 3.7.
Notice that all the constraints needed in the proof of the lower bound and in the proof of the upper bound are ultimately a matter of choosing $\rho_{i,2}$ large enough; this guarantees that we may get the $N_{i}$ obtained at the end of the two proofs to be the same integer.
Appendix A Auxiliary results
We start by proving the technical lemma at the end of Subsection 3.1.
Proof of Lemma 3.5.
Figure 4: The auxiliary tools we constructed.
We need some auxiliary geometric tools to prove Lemma 3.5, and for the sake of notation, we omit $C$ in the subscript. Without loss of generality, we let $\boldsymbol{\Gamma}\colon[-1,1)\to\mathbb{R}^{2}$ be the counter-clockwise arc-length parametrisation of $\partial C$ such that $\boldsymbol{\Gamma}(0)=\mathbf{0}$, and we write $\boldsymbol{\Gamma}^{\prime}$ as its velocity (in particular, $\left|\boldsymbol{\Gamma}^{\prime}\right|\equiv 1$). For a chord $K(\theta,\lambda)$, we define $a(\theta,\lambda)$ and $b(\theta,\lambda)$ to be the parametrisation by $\boldsymbol{\Gamma}$ of the extreme points of $K(\theta,\lambda)$, with the convention that
$\boldsymbol{\Gamma}(a(\theta,\lambda))-\boldsymbol{\Gamma}(b(\theta,\lambda))=|K(\theta,\lambda)|\mathbf{u}^{\prime}(\theta).$
Moreover, we set $p(\theta)$ to be the parametrisation by $\boldsymbol{\Gamma}$ of the point at $\partial C$ whose inner normal is $\mathbf{u}(\theta)$. Finally, we split the chord $K(\theta,\lambda)$ into two semi-chords $K^{+}(\theta,\lambda)$ and $K^{-}(\theta,\lambda)$ as displayed in Figure 4. Now, choose $\theta_{0}$ and $\lambda_{0}$ so that, for $\lambda\in[0,\lambda_{0})$ and $\theta\in[\pi/2,\theta_{0})$, it holds $\boldsymbol{\Gamma}_{2}^{\prime}(b(\theta,\lambda))\leq 1/10$. The latter requirement bounds the endpoints of the chords $K(\theta,\lambda)$ to the curve segment constructed in Subsection 3.1. Thus, we may conclude once we prove that, for $\lambda\in[0,\lambda_{0})$ and $\pi/2\leq\theta_{1}<\theta_{2}<\theta_{0}$, it holds
$\left|K^{+}(\theta_{2},\lambda)\right|\geq\left|K^{+}(\theta_{1},\lambda)\right|\geq\left|K^{-}(\theta_{1},\lambda)\right|.$
We only show the inequality $\left|K^{+}(\theta_{2},\lambda)\right|\geq\left|K^{+}(\theta_{1},\lambda)\right|$, since the argument for the other one is analogous. We begin by noting that, for $i\in\{1,2\}$, it holds
$\lambda=\int_{0}^{b(\theta_{i},\lambda)-p(\theta_{i})}\mathbf{u}(\theta_{i})\cdot\boldsymbol{\Gamma}^{\prime}(p(\theta_{i})+s)\,\mathrm{d}s$ (A.1)
and
$\left|K^{+}(\theta_{i},\lambda)\right|=\int_{0}^{b(\theta_{i},\lambda)-p(\theta_{i})}\mathbf{u}(\theta_{i}-\pi/2)\cdot\boldsymbol{\Gamma}^{\prime}(p(\theta_{i})+s)\,\mathrm{d}s.$ (A.2)
Since the derivative of the argument of $\boldsymbol{\Gamma}^{\prime}$ is exactly the signed curvature $\kappa$, then for $i\in\{1,2\}$ we obtain
$\boldsymbol{\Gamma}^{\prime}(p(\theta_{i}))=\mathbf{u}(\theta_{i}-\pi/2)\quad\text{and}\quad\boldsymbol{\Gamma}^{\prime}(p(\theta_{i})+s)=\mathbf{u}\left(\theta_{i}-\pi/2+\int_{0}^{s}\kappa(p(\theta_{i})+t)\,\mathrm{d}t\right),$
so that, by some basic trigonometry, we may rewrite (A.1) and (A.2) as
$\lambda=\int_{0}^{b(\theta_{i},\lambda)-p(\theta_{i})}\sin\left(\int_{0}^{s}\kappa(p(\theta_{i})+t)\,\mathrm{d}t\right)\,\mathrm{d}s$ (A.3)
and
$\left|K^{+}(\theta_{i},\lambda)\right|=\int_{0}^{b(\theta_{i},\lambda)-p(\theta_{i})}\cos\left(\int_{0}^{s}\kappa(p(\theta_{i})+t)\,\mathrm{d}t\right)\,\mathrm{d}s.$ (A.4)
Notice that, by our construction of the curve segment $P$ in Subsection 3.1, it holds the inequality
$\kappa(p(\theta_{1})+t)\geq\kappa(p(\theta_{2})+t)\quad\text{for}\quad t\in\left[0,b(\theta_{0},\lambda_{0})-p(\theta_{2})\right),$
but this, combined with (A.3) and the monotonicity of the sine function in $0,\pi/2)$, implies that
$b(\theta_{1},\lambda)-p(\theta_{1})\leq b(\theta_{2},\lambda)-p(\theta_{2}),$
and in turn, combining the latter with (A.4) and the monotonicity of the cosine function in $0,\pi/2)$, implies
$\left|K^{+}(\theta_{2},\lambda)\right|\geq\left|K^{+}(\theta_{1},\lambda)\right|.$
Before proceeding with the proof of Proposition 3.3, we first state the following auxiliary result. We will not present its proof here, as it is a simple exercise; however, the reader may find a proof in [BER25, Lemma 8.1].
Lemma A.1.
Let $\alpha$ and $\beta$ be positive numbers, and let $g\colon[0,+\infty)\to[0,+\infty)$ be such that
$g(x)\approx\begin{cases}x^{\alpha}&\textnormal{if}\quad 0\leq x<1\\ x^{\beta}&\textnormal{if}\quad x\geq 1\end{cases}.$
If $x_{y}$ is such that $g(x_{y})=y$, then it holds
$x_{y}\approx\begin{cases}y^{1/\alpha}&\textnormal{if}\quad 0\leq y<1\\ y^{1/\beta}&\textnormal{if}\quad y\geq 1\end{cases}.$
[…]
$\left|K^{+}_{A(\beta)}(\theta+\pi/2,\lambda)\right|\geq\left|K^{-}_{A(\beta)}(\theta+\pi/2,\lambda)\right|.$
[…]
Proof of Proposition 3.3.
[…]
$\left|K_{A(\beta)}^{+}(\theta+\pi/2,\lambda)\right|\approx x_{+}(\theta+\pi/2,\lambda)-x_{o}(\theta+\pi/2).$
For the sake of brevity, we will write $x_{o}=x_{o}(\theta+\pi/2)$ and $x_{+}=x_{+}(\theta+\pi/2,\lambda)$. Now, notice that $x_{+}$ is the abscissa of the (right) intersection of the curve $y=|x|^{\beta}$ with the straight line
$y=(x-x_{o})\beta x_{o}^{\beta-1}+x_{o}^{\beta}+\frac{\lambda}{\cos\theta}.$
Equating, and with the normalisation $z=\frac{x-x_{o}}{x_{o}}$, we get to the equation
$f(z)=|z+1|^{\beta}-z\beta-1=\frac{\lambda}{x_{o}^{\beta}\cos\theta},$ (A.5)
[…]
$f(z)=\beta(\beta-1)\int_{0}^{z}(1+t)^{\beta-2}(z-t)\,\mathrm{d}t.$
Notice that for $z\in[0,1)$ it holds
$\int_{0}^{z}(1+t)^{\beta-2}(z-t)\,\mathrm{d}t\approx\int_{0}^{z}(z-t)\,\mathrm{d}t\approx z^{2}.$
On the other hand, for $z\in[1,+\infty)$ it holds
$\begin{split}\int_{0}^{z}(1+t)^{\beta-2}(z-t)\,\mathrm{d}t&=\int_{0}^{z/2}(1+t)^{\beta-2}(z-t)\,\mathrm{d}t+\int_{\frac{z}{2}}^{z}(1+t)^{\beta-2}(z-t)\,\mathrm{d}t\\ &\approx z\int_{0}^{z/2}(1+t)^{\beta-2}\,\mathrm{d}t+z^{\beta-2}\int_{\frac{z}{2}}^{z}(z-t)\,\mathrm{d}t\\ &=\frac{z}{\beta-1}\left(\left(1+z/2\right)^{\beta-1}-1\right)+z^{\beta-2}\frac{z^{2}}{8}\approx_{\beta}z^{\beta}.\end{split}$
Hence, we get
$f(z)\approx_{\beta}\begin{cases}z^{2}&\text{if}\quad 0\leq z<1\\ z^{\beta}&\text{if}\quad z\geq 1\end{cases},$
[…]
$\frac{x_{+}-x_{o}}{x_{o}}\approx_{\beta}\begin{cases}\lambda^{1/2}x_{o}^{-\beta/2}&\text{if}\quad 0\leq\lambda x_{o}^{-\beta}<1\\ \lambda^{1/\beta}x_{o}^{-1}&\text{if}\quad\lambda x_{o}^{-\beta}\geq 1\end{cases}.$ (A.6)
Finally, by the definition of $x_{o}$, we have
$\beta x_{o}^{\beta-1}=\left.\frac{d}{dx}x^{\beta}\right|_{x=x_{o}}=\tan\theta,$
[…]
$x_{o}\approx\theta^{\frac{1}{\beta-1}}.$
The conclusion follows by a rearrangement of the terms in (A.6). ∎
Acknowledgements
I am grateful to my advisors, Luca Brandolini, Leonardo Colzani, Giacomo Gigante, and Giancarlo Travaglini, for their support and all the valuable discussions.